# Handling Reward Misspecification in the Presence of Expectation Mismatch

### v1

@@ 1 Introduction
Given the accelerating pace of advancement within AI, creating agents that can detect and handle incorrectly specified objectives has become an evermore pressing problem Dafoe et al. (2020). This has resulted in the development of several methods for addressing, among other forms of objective functions, misspecified rewards (cf. Hadfield-Menell et al. (2017)). Unfortunately, these works operate under an implicit definition of what constitutes an incorrectly specified reward function. The few works that have tried to formalize how an agent could satisfy human objectives do so by avoiding the objective specification step altogether Hadfield-Menell et al. (2016).
Our primary objective with this paper is to start with an explicit formalization of what constitutes a misspecified reward function and the potential reasons why the user may have provided one in the first place. However, our motivation for providing such a formalization goes beyond just the pedagogical. We see that such a formalization provides us with multiple insights of great practical importance.
Our primary formulation will try to capture the basic intuition that when a user proposes a reward function or any objective specification for that matter, they have a specific outcome [^1] In this paper, we will focus on cases where expected/desired outcomes correspond to the agent achieving certain states. Section 3 provides a formal definition. (or a set of acceptable outcomes) that they hope the agent will accomplish. The reward function proposed by the user is one whose maximization, they believe, will result in a policy that will achieve the desired outcome. Figure 1 shows how users derive their reward specifications. As presented, their reward function will be informed by their beliefs about the agent and its capabilities (i.e., the user’s theory of mind Apperly and Butterfill (2009) of the system) and their ability to reason about how the agent will act given a specific reward function.
A given reward function is said to be misspecified when the policy identified by the agent doesn’t satisfy the outcomes the user was expecting. As evident from previous discussions, the reasons for such misspecification could include users’ incorrect beliefs about the agent and their limited inferential capabilities.
We will formalize this intuition about reward specification and, by extension, that of misspecification under the more general framework of Expectation Alignment (EAL). EAL framework will allow us to show that there are scenarios under which there doesn’t exist a reward function that will result in the agent satisfying the human expectation. This invalidates all approaches that try to identify a ‘true’ human reward function, which is then passed onto the agent. Secondly, we will see how we can use EAL to develop novel algorithms for handling reward misspecification that explicitly leverages potential causes of misalignment, to come up with more effective queries to identify the human’s original intent.
To summarize, the contributions of the paper are:
@@
  • We develop a novel query-based algorithm to solve a specific instance of EAL problems (obtained by fixing certain parameters of the problems).
@@ 2 Background
We will primarily focus on problems that can be expressed as a Markov Decision Process or an MDP Puterman (2014). An MDP is usually represented by a tuple of the form $\mathcal{M}=\langle S,A,T,R,\gamma,s_{0}\rangle$. Under this notation scheme, $S$ captures the set of states, $A$ the set of actions, $T:S\times A\times S\rightarrow[0,1]$ is the transition function such that $T(s,a,s^{\prime})$ captures the probability of an action $a$ causing the state to transition from $s$ to $s^{\prime}$, $R:S\rightarrow\mathbb{R}$ is the reward function, $\gamma\in[0,1)$ the discount factor and finally $s_{0}\in S$ is the initial state (the agent is expected to start from $s_{0}$). Since we are focusing on problems related to reward specification, we will separate out all non-reward components of an MDP model, refer to it as the domain, and denote it as $\mathcal{D}=\langle S,A,T,\gamma,s_{0}\rangle$.
@@
In this paper, we will be heavily utilizing the notion of discounted occupancy frequency or just occupancy frequency, $x^{\pi}:S\times A\rightarrow[0,1]$ Poupart (2005), associated with a policy $\pi$. There are multiple ways to interpret occupancy frequency, but one of the most common ones is to think of them as the steady state probability associated with a policy $\pi$. In particular, the occupation frequency $x^{\pi}(s,a)$ captures the frequency with which an action $a$ would be executed in a state $s$ under the policy $\pi$. Sometimes, we will also need to identify the frequency with which a policy $\pi$ would visit a state $s$. We can obtain this frequency by summing over the values of $x^{\pi}(s,a)$ for all actions i.e. $x^{\pi}(s)=\sum_{a}x^{\pi}(s,a)$.
It is possible to reformulate the value obtained under a policy in terms of its occupancy frequency and rewards. One can also identify the optimal value and, thereby, the optimal policy by solving the following linear program (LP) expressed using occupancy frequency variables:
$\begin{split}\max_{x}\sum_{s,a}x(s,a)\times r(s)\\[-8.0pt] \text{s.t.}\quad\forall s\in S,\sum_{a}x(s,a)=\delta(s,s_{0})+\\[-8.0pt] \gamma\times\sum_{s^{\prime},a^{\prime}}x(s^{\prime},a^{\prime})\times T(s,a,s^{\prime})\end{split}$ (1)
@@ 3 Expectation Alignment Framework
To build such a forward model, we need a formal method to represent behavioral expectations in the context of MDPs. While several promising mathematical formalisms could be adopted to represent a human’s behavioral expectations, we will focus on using the notion of occupancy frequency. There are multiple reasons for this choice. For one, this is a natural generalization of the notions of reachability. Psychological evidence supports that people use the notion of goals in decision-making Simon (1977). While goals relate to the idea of reaching desirable end states, the notion of occupancy frequency allows us to extend it to intermediate ones. It even allows us to express probabilistic notions of reachability. Secondly, occupancy frequencies present a very general notion of problems that can be expressed as MDPs. As discussed in Section 2, the value of any given policy can be represented in terms of their occupancy frequency Poupart (2005). Finally, the popular use-cases in the space of reward misspecification can be captured using occupancy frequency (See Section 5).
@@
For cases where the human wants the robot to completely avoid a particular state $s$, there would be a corresponding expectation element $\langle\{s\},=,0\rangle$; on the other hand, if the human wants the robot to visit certain states, then there will be expectation elements that try to enforce high occupancy frequencies for those states. In the simplest settings, humans might convert such expectations to rewards by setting low or even negative rewards to the states that need to be avoided and high rewards to the states they want the robot to visit. One question that is worth considering is why humans don’t just communicate this expectation set. Unfortunately, communicating the entire set may be relatively inefficient compared to coming up with a reward function (for example, one might be able to compactly represent the reward function as a piece of code, while this set may not). This could also be the case if the number of states is very high or even infinite, but even in these settings, the reward function could still be specified in terms of features. Secondly, even if they choose to specify desirable or undesirable states, the expectation set may contain states that the human incorrectly thinks are impossible (for example, the robot jumping onto the table) and may choose to ignore them. This is also related to the problem of unspecified side-effects Zhang et al. (2018).
It is worth noting that Definition 1 merely provides a mathematical formulation that is general enough to capture human expectations. We don’t expect people to maintain an explicit set of numeric constraints on occupancy frequencies in their minds. Regardless of what form their actual expectation takes, as long as it can be expressed as some ordering on how the states should be visited, it can be captured using sets of the form provided in the above definition. One could, in theory, also capture non-Markovian expectations using such formalisms. For example, there may be cases where the user might want the robot to visit a set of states in a specific sequence. We can capture such cases by creating augmented states that can track whether or not the robot visited the states in sequence. This method also allows Markovian reward functions to capture such considerations (cf. Abel et al. (2021)), which is not surprising since the expressivity of occupancy frequency in many ways parallels that of Markovian reward functions.
We will claim that a policy satisfies an expectation element for a domain if the corresponding occupancy frequency relation holds for that policy in that domain, or more formally:
@@
To relate these expectations to the reward specification process, we first need a way to represent the human decision-making process. For this discussion, we will start with a very abstract model of human decision-making, which we will later ground in Section 4. In particular, we will represent human decision-making as constituting a planning function $\mathcal{P}^{H}$, that maps a given model to a (possibly empty) set of policies.
@@
For a given model $\mathcal{M}$, the set of policies that the human thinks will be selected is given by the planning function $\mathcal{P}^{H}:\mathbb{M}\rightarrow\Pi$, where $\mathbb{M}$ is the space of possible models and $\Pi$ the space of possible policies.
It is worth noticing that, as with $\mathcal{D}^{H}$, $\mathcal{P}^{H}$ is a reflection of the human’s belief about the robot’s planning capabilities. In many cases, it is possible for the human to ascribe a lower level of cognitive capability to the robot. However, for this paper, we will ignore such considerations.
With these definitions in place, we can describe how humans would choose an acceptable reward specification. Specifically, the human would think a reward function is sufficient if all policies they believe could be selected under this specification will satisfy their expectations, or more formally
@@
In this definition, we can already see the outlines of the various sources of misspecification. For one, humans use their understanding of the task and the robot’s capabilities (captured by $\mathcal{D}^{H}$) to choose the reward specification. The true robot capabilities or task-level constraints could drastically differ from human beliefs. Next, the policies they anticipated being selected ($\mathcal{P}^{H}(\langle\mathcal{D}^{H},\mathcal{R}\rangle)$) could be very different from what might be chosen by the robot. There is also the additional possibility that the human is not even able to correctly evaluate whether an expectation element holds for a given policy, especially given the fact that people are known to be notoriously bad at handling probabilities Tversky and Kahneman (1983); Kahneman and Tversky (1981). In many cases, the expectations might be limited to the occupancy frequencies being forced to be zero (avoid certain states) or taking a non-zero value (should try to visit certain states). In such cases, the human could ignore the probabilities and just reason using a determinized version of the model Yoon et al. (2008). Even here, there is a possibility that the human could overlook some improbable but possible paths that might cause the expectation element to be unsatisfied.
More explicitly, we will deem a reward function to be misspecified if there is at least one robot optimal policy where at least one expectation set element is unsatisfied.
@@
We can prove this by constructing an example where it holds. Consider a case where, in the set of states the human wants the agent to visit (i.e., the sum of occupancy frequency $>0$), one might need to reward some states in the human model and others in the robot model. This might be because, in the human model, visiting the other set might result in the robot slipping into states that shouldn’t be visited, and the reverse holds in the robot model. ∎
As such, one can’t talk about a single ‘true’ reward function that holds for both domains. This rules out the possibility of using methods like inverse reward design Hadfield-Menell et al. (2017), which aims to generate a single reward function that applies to both the intended and actual environments. On the other hand, the underlying expectations are directly transferrable across the domains.
This brings us to what the robot’s objective should be, namely, to come up with an expectation-aligned policy, i.e., one that satisfies human expectations in the robot task domain, when possible, or recognize when it cannot do so.
@@
The primary challenge to generating expectation-aligned policies is that the expectation set is not directly available. However, unlike previous works in this space, the fact that we have a model that maps the expectations to the reward function means that there is an opportunity to develop a more diverse set of possible solution approaches. For one, we could now look at the possibility of starting by learning the human model (using methods similar to the ones described in Gong and Zhang (2020); Reddy et al. (2018)) and use that information to produce estimates of $\mathbb{E}^{H}$ from the reward specification. Along the same lines, we could leverage psychologically feasible models as a stand-in for $\mathcal{P}^{H}$ to further improve the accuracy of our estimates (possible candidates include noisy-rational model Jeon et al. (2020)). In the next section, we show how we can develop effective algorithms for generating expectation-aligned policies by looking at an instance where the expectation set is limited to a certain form.
@@ 4 Identifying Expectation-Aligned Policies
We will ground the proposed framework in a use case with a relatively simple planning function $\mathcal{P}$ that returns the set of optimal policies, and where the expectation set takes a particular form. Expressly, we assume that there is an unknown set of states in the expectation set that cannot be visited, i.e., there exists some $e\in\mathbb{E}^{H}$, takes the form $\langle\{s\},=,0\rangle$ and some states that they would like the robot to visit (i.e., $e$ of the form $\langle\{s\},>,0\rangle$). We will call the states that are part of the first set of expectations the forbidden state set $\mathcal{S}^{\mathcal{F}}$ and the second one the goal set $\mathcal{S}^{\mathcal{G}}$. The expectation set described here corresponds to the widely studied reward misspecification problem type, where the robot needs to avoid negative side effects while achieving some goal (cf. Saisubramanian et al. (2022)).
@@
There exists no state $s\in\mathcal{S}^{\mathcal{F}}$ and policy $\pi\in\Pi^{*}_{\mathcal{M}^{H}}$, such that the condition $x^{\pi}(s)>0$ is true.
Next, we can also see that the states part of $\mathcal{S}^{\mathcal{G}}$ must be reachable under every optimal policy.
@@
For every state $s\in\mathcal{S}^{\mathcal{G}}$ and policy $\pi\in\Pi^{*}_{\mathcal{M}^{H}}$ the condition $x^{\pi}(s)>0$ must always be true.
One might be tempted to generate the sets $\mathcal{S}^{\mathcal{F}}$ and $\mathcal{S}^{\mathcal{G}}$ by looking at all states with low and high reward values, respectively. However, there are multiple problems with such a naive approach. Firstly, it is well-known that people encode solution strategies into the reward function in addition to eventual goals Booth et al. (2023). This means there may be high-reward value states that are not part of $\mathcal{S}^{\mathcal{G}}$ or relatively low-reward states that are not part of $\mathcal{S}^{\mathcal{F}}$. These may have been added to get the robot to behave in certain ways. Secondly, humans may ignore states they think are guaranteed to be reached or will never be reached. In our running example, human ignores the walkway because they think the robot is incapable of using it.
While we can’t exactly calculate $\mathcal{S}^{\mathcal{F}}$ and $\mathcal{S}^{\mathcal{G}}$, we will proceed to show how we can generate supersets of these sets. Specifically, we can calculate the set of all states that aren’t reachable in any optimal policies (denoted as $\widehat{\mathcal{S}}^{\mathcal{F}}\supseteq\mathcal{S}^{\mathcal{F}}$) and the set of all states that are reachable in every optimal policy ($\widehat{\mathcal{S}}^{\mathcal{G}}\supseteq\mathcal{S}^{\mathcal{G}}$).
In particular, we can generate these sets using modified forms of the LP formulation discussed in Equation 1. To test whether a given state $s_{i}$ is part of $\widehat{\mathcal{S}}^{\mathcal{F}}$ or not, we will look at an LP that will try to identify an optimal policy that will visit $s_{i}$. Specifically, the LP would look like:
$\begin{split}\max_{x}\sum_{s,a}x(s,a)\times r(s)+\alpha\times(\sum_{a}x(s_{i},a))\\[-8.0pt] \text{s.t.}\quad\forall s\in S,\sum_{a}x(s,a)=\delta(s,s_{0})+\\[-8.0pt] \gamma\times\sum_{s^{\prime},a^{\prime}}x(s^{\prime},a^{\prime})\times T^{H}(s,a,s^{\prime})\\[-8.0pt] \sum_{a}x(s_{0},a)=V^{*}_{s_{0}}\\[-8.0pt] \end{split}$ (2)
Where $\alpha$ is some positive valued coefficient and $V^{*}_{s_{0}}$ is the value of the optimal policy in the starting state $s_{0}$. We can calculate $V^{*}_{s_{0}}$ by solving Equation 1 on the human model with the specified reward.
By adding the new term to the objective, we provide higher objective value to policies that visit the state $s_{i}$. At the same time, the constraint $\sum_{a}x(s_{0},a)=V^{*}_{s_{0}}$ ensures we only consider optimal policies. We can now formally state the following property for the above LP formulation.

### v2

# Expectation Alignment: Handling Reward Misspecification in the Presence of Expectation Mismatch

@@ 1 Introduction
Given the accelerating pace of advancement within AI, creating agents that can detect and handle incorrectly specified objectives has become an evermore pressing problem [Dafoe et al., 2020]. This has resulted in the development of several methods for addressing, among other forms of objective functions, misspecified rewards (cf. Hadfield-Menell et al. [2017]). Unfortunately, these works operate under an implicit definition of what constitutes an incorrectly specified reward function. The few works that have tried to formalize how an agent could satisfy human objectives do so by avoiding the objective specification step altogether [Hadfield-Menell et al., 2016].
Our primary objective with this paper is to start with an explicit formalization of what constitutes a misspecified reward function and the potential reasons why the user may have provided one in the first place. However, our motivation for providing such a formalization goes beyond just the pedagogical. We see that such a formalization provides us with multiple insights of practical importance.
Our formulation will follow the intuition set by recent work (cf. [Abel et al., 2021]) that looks at reward functions as a means of specifying a task as opposed to being an end to itself. As such, the user starts with a target outcome or behavior or a set of outcomes or behaviors  [^1] In this paper, we will focus on cases where expected/desired outcomes correspond to the agent achieving certain states. Section 3 provides a formal definition. that they want the agent to achieve. The reward function proposed by the user is one whose maximization, they believe, will result in a policy that will achieve the desired outcome. Figure 1 shows how users derive their reward specifications. As presented, their reward function will be informed by their beliefs about the agent and its capabilities (i.e., the user’s theory of mind [Apperly and Butterfill, 2009] about the agent) and their ability to reason about how the agent will act given a specific reward function.
@@
A given reward function is said to be misspecified when the policy identified by the agent for the reward function doesn’t satisfy the user’s expectations. As discussed, the reasons for such misspecification could include users’ incorrect beliefs about the agent and their limited inferential capabilities [Sreedharan, 2022]. To see how such misspecification might occur, consider a simple planning problem that can be captured by a Markov Decision Process (MDP) with only three states. Consider a problem where an agent can only perform two possible actions. These correspond to pressing two different switches. Each switch takes the agent to one of the two possible end states where the agent can no longer act. One corresponds to a safe end state, while the other corresponds to an unsafe one. Let’s consider a human supervisor who wants the agent to go to the safe end state and, as such, identifies a reward function that they believe will result in such a state. If the supervisor is confused about which button leads to which state, they may end up coming up with a reward function that will result in the exact opposite outcome. Please note that the problem here isn’t that the agent can’t achieve the desired outcome but rather that optimizing for the specified reward will not result in the desired outcome. In fact, given the human’s confusion about the buttons, no reward function exists that could lead the agent to the desired outcome, which the human will agree is a correct reward function for the task.
We will formalize this intuition about reward specification and, by extension, that of misspecification under the more general framework of Expectation Alignment (EAL). EAL framework will allow us to capture scenarios, like the one discussed above, where the human cannot come up with a reward that satisfies the expectation in both their and the agent’s models. This invalidates all approaches that try to identify a ‘true’ human reward function, which is then passed onto the agent. Secondly, we will see how we can use EAL to develop novel algorithms for handling reward misspecification that explicitly leverages potential causes of misalignment to come up with more effective queries to identify the human’s original intent. To summarize, the contributions of the paper are:
@@
  • We develop a novel query-based algorithm to solve a specific instance of EAL problems.
@@
In the related work section (Section 5), we will also show how existing works on handling objective misspecification relate to our proposed framework. Since it will leverage the specifics of our proposed framework, we will look at the related work after defining our basic framework.
@@ 2 Background
We will primarily focus on problems that can be expressed as a Markov Decision Process or an MDP [Puterman, 2014]. An MDP is usually represented by a tuple of the form $\mathcal{M}=\langle S,A,T,R,\gamma,s_{0}\rangle$ [^2] We can additionally use control constraints to limit the actions possible from each state, as in the example in the introduction. However, we will leave it out of the formulation to support a more concise formulation.. Under this notation scheme, $S$ captures the set of states, $A$ the set of actions, $T:S\times A\times S\rightarrow[0,1]$ is the transition function such that $T(s,a,s^{\prime})$ captures the probability of an action $a$ causing the state to transition from $s$ to $s^{\prime}$, $R:S\times A\rightarrow\mathbb{R}$ is the reward function, $\gamma\in[0,1)$ the discount factor and finally $s_{0}\in S$ is the initial state (the agent is expected to start from $s_{0}$). Since we are focusing on problems related to reward specification, we will separate out all non-reward components of an MDP model, refer to it as the domain, and denote it as $\mathcal{D}=\langle S,A,T,\gamma,s_{0}\rangle$.
@@
In this paper, we will be heavily utilizing the notion of discounted occupancy frequency or just occupancy frequency, $x^{\pi}:S\times A\rightarrow[0,1]$ [Poupart, 2005], associated with a policy $\pi$. There are multiple ways to interpret occupancy frequency, but one of the most common ones is to think of them as the steady state probability associated with a policy $\pi$. In particular, the occupation frequency $x^{\pi}(s,a)$ captures the frequency with which an action $a$ would be executed in a state $s$ under the policy $\pi$. Sometimes, we will also need to identify the frequency with which a policy $\pi$ would visit a state $s$. We can obtain this frequency by summing over the values of $x^{\pi}(s,a)$ for all actions, i.e., $x^{\pi}(s)=\sum_{a}x^{\pi}(s,a)$. It is possible to reformulate the value obtained under a policy in terms of its occupancy frequency and rewards. One can also identify the optimal value and, thereby, the optimal policy by solving the following linear program (LP) expressed using occupancy frequency variables:
$\begin{split}\max_{x}\sum_{s,a}x(s,a)\times r(s,a)\\[-5.0pt] \text{s.t.}\quad\forall s\in S,\sum_{a}x(s,a)=\delta(s,s_{0})+\gamma\times\sum_{s^{\prime},a^{\prime}}x(s^{\prime},a^{\prime})\times T(s^{\prime},a^{\prime},s)\end{split}$ (1)
@@ 3 Expectation Alignment Framework
To build such a forward model, we need a formal method to represent behavioral expectations in the context of MDPs. While several promising mathematical formalisms could be adopted to represent a human’s behavioral expectations, we will focus on using the notion of occupancy frequency. There are multiple reasons for this choice. For one, this is a natural generalization of the notions of reachability. Psychological evidence supports that people use the notion of goals in decision-making [Simon, 1977]. While goals relate to the idea of reaching desirable end states, the notion of occupancy frequency allows us to extend it to intermediate ones. It even allows us to express probabilistic notions of reachability. Secondly, occupancy frequencies present a very general notion of problems that can be expressed as MDPs. As discussed in Section 2, the value of any given policy can be represented in terms of their occupancy frequency [Poupart, 2005]. Finally, the popular use cases in the space of reward misspecification can be captured using occupancy frequency (See Section 5).
@@
For cases where the human wants the robot to completely avoid a particular state $s$, there would be a corresponding expectation element $\langle\{s\},=,0\rangle$; on the other hand, if the human wants the robot to visit certain states, then there will be expectation elements that try to enforce high occupancy frequencies for those states. In the simplest settings, humans might convert such expectations to rewards by setting low or even negative rewards to the states that need to be avoided and high rewards to the states they want the robot to visit. One question that is worth considering is why humans don’t just communicate this expectation set. Unfortunately, communicating the entire set may be relatively inefficient compared to coming up with a reward function (for example, one might be able to compactly represent the reward function as a piece of code, while this set may not). This could also be the case if the number of states is very high or even infinite, but even in these settings, the reward function could still be specified in terms of features. Secondly, even if they choose to specify desirable or undesirable states, the expectation set may contain states that the human incorrectly thinks are impossible (for example, the robot jumping onto the table) and may choose to ignore them. This is also related to the problem of unspecified side-effects [Zhang et al., 2018].
It is worth noting that Definition 1 merely provides a mathematical formulation that is general enough to capture human expectations. We don’t expect people to maintain an explicit set of numeric constraints on occupancy frequencies in their minds. Regardless of what form their actual expectation takes, as long as it can be expressed as some ordering on how the states should be visited, it can be captured using sets of the form provided in the above definition. One could, in theory, also capture non-Markovian expectations using such formalisms. For example, there may be cases where the user might want the robot to visit a set of states in a specific sequence. We can capture such cases by creating augmented states that can track whether or not the robot visited the states in sequence. This method also allows Markovian reward functions to capture such considerations (cf. [Abel et al., 2021]), which is unsurprising since the expressivity of occupancy frequency parallels that of Markovian reward functions.
We will claim that a policy satisfies an expectation element for a domain if the corresponding occupancy frequency relation holds for the policy in that domain, or more formally:
@@
To relate these expectations to the reward specification process, we first need a way to represent the human decision-making process. For this discussion, we will start with a very abstract model of human decision-making, which we will later ground in Section 4. In particular, we will represent human decision-making using a planning function $\mathcal{P}^{H}$, that maps a given model to a (possibly empty) set of policies.
@@
For a given model $\mathcal{M}$, the set of policies that the human thinks will be selected is given by the planning function $\mathcal{P}^{H}:\mathbb{M}\rightarrow 2^{\Pi}$, where $\mathbb{M}$ is the space of possible models and $\Pi$ the space of possible policies.
It is worth noticing that, as with $\mathcal{D}^{H}$, $\mathcal{P}^{H}$ is a reflection of the human’s belief about the robot’s model and planning capabilities. In many cases, humans might ascribe a lower level of cognitive capability to the robot. However, for this paper, we will ignore such considerations. With these definitions in place, we can describe how humans would choose an acceptable reward specification. Specifically, the human would think a reward function is sufficient if all policies they believe could be selected under this specification will satisfy their expectations, or more formally
@@
In this definition, we can already see the outlines of the various sources of misspecification. For one, humans use their understanding of the task and the robot’s capabilities (captured by $\mathcal{D}^{H}$) to choose the reward specification. The true robot capabilities or task-level constraints could drastically differ from human beliefs. Next, the policies they anticipated being selected ($\mathcal{P}^{H}(\langle\mathcal{D}^{H},\mathcal{R}\rangle)$) could be very different from what might be chosen by the robot. There is also the additional possibility that the human is not even able to correctly evaluate whether an expectation element holds for a given policy, especially given the fact that people are known to be notoriously bad at handling probabilities [Tversky and Kahneman, 1983, Kahneman and Tversky, 1981]. In many cases, the expectations might be limited to the occupancy frequencies being forced to be zero (avoid certain states) or taking a non-zero value (should try to visit certain states). Here, the human could ignore the probabilities and reason using a determinized version of the model [Yoon et al., 2008]. Even here, there is a possibility that humans could overlook some paths that might cause the expectation element to be unsatisfied.
In general, we will deem a reward function misspecified if at least one robot optimal policy has at least one unsatisfied expectation set element.
@@
We can prove this by constructing an example where it holds. Consider the example discussed in the introduction. Here, there is a state that the human wants the robot to visit (i.e., the occupancy frequency $>0$), and there is a state that the human wants to avoid(i.e., the occupancy frequency $=0$). To achieve the expected behavior in the human model, they have to choose a reward function that will result in an optimal policy where the button they believe will lead to the goal state has to be pressed. However, if the functionality of the switches is reversed in the robot model, none of those rewards will result in a robot policy that will lead to the state that the human wanted to visit. This shows that, for this example, there exists no reward function that will result in a policy that satisfies the expectation set. Thus proving the theorem statement. ∎
This theorem shows that one can’t talk about a single ‘true’ reward function that holds for both domains. In the example, as far as the human is concerned, the reward they came up with is the true reward since it results in the exact behavior they wanted. In fact, any reward function that results in the robot achieving its goal will be considered incorrect since it will never lead to the expected behavior in their model. This rules out the possibility of using methods like inverse reward design [Hadfield-Menell et al., 2017], which aims to generate a single reward function that applies to both the intended and actual environments. On the other hand, the underlying expectations are directly transferrable across the domains. This brings us to what the robot’s objective should be, namely, to come up with an expectation-aligned policy, i.e., one that satisfies human expectations in the robot task domain, when possible, or recognize when it cannot do so.
@@
The primary challenge to generating expectation-aligned policies is that the expectation set is not directly available. However, unlike previous works in this space, the fact that we have a model that maps the expectations to the reward function means that there is an opportunity to develop a more diverse set of possible solution approaches. For one, we could now look at the possibility of starting by learning the human model (using methods similar to the ones described by Gong and Zhang [2020] and Reddy et al. [2018]) and use that information to produce estimates of $\mathbb{E}^{H}$ from the reward specification. Along the same lines, we could leverage psychologically feasible models as a stand-in for $\mathcal{P}^{H}$ to further improve the accuracy of our estimates (possible candidates include noisy-rational model [Jeon et al., 2020]). In the next section, we show how we can develop effective algorithms for generating expectation-aligned policies by looking at an instance where the expectation set is limited to a certain form.
@@ 4 Identifying Expectation-Aligned Policies
We will ground the proposed framework in a use case with a relatively simple planning function $\mathcal{P}$ that returns the set of optimal policies and where the expectation set takes a particular form. Specifically, we assume that there is an unknown set of states in the expectation set that cannot be visited, i.e., there exists some $e\in\mathbb{E}^{H}$, takes the form $\langle\{s\},=,0\rangle$ and some states that they would like the robot to visit (i.e., $e$ of the form $\langle\{s\},>,0\rangle$). We will call the states that are part of the first set of expectations the forbidden state set $\mathcal{S}^{\mathcal{F}}$ and the second one the goal set $\mathcal{S}^{\mathcal{G}}$. The expectation set described here corresponds to the widely studied reward misspecification problem type, where the robot needs to avoid negative side effects while achieving the goal (cf. Saisubramanian et al. [2022]).
@@
There exists no state $s\in\mathcal{S}^{\mathcal{F}}$ and policy $\pi\in\Pi^{*}_{\mathcal{M}^{H}}$, such that $x^{\pi}(s)>0$ is true.
The proof sketch for the proposition is provided in Appendix 8.1. Next, we can also see that the states part of $\mathcal{S}^{\mathcal{G}}$ must be reachable under every optimal policy.
@@
For every state $s\in\mathcal{S}^{\mathcal{G}}$ and policy $\pi\in\Pi^{*}_{\mathcal{M}^{H}}$, $x^{\pi}(s)>0$ must always be true.
The proof sketch for the above proposition is again provided in Appendix 8.1. One might be tempted to generate the sets $\mathcal{S}^{\mathcal{F}}$ and $\mathcal{S}^{\mathcal{G}}$ by looking at all states with low and high reward values, respectively. However, there are multiple problems with such a naive approach. Firstly, it is well-known that people encode solution strategies into the reward function in addition to eventual goals [Booth et al., 2023]. This means there may be high-reward value states that are not part of $\mathcal{S}^{\mathcal{G}}$ or relatively low-reward states that are not part of $\mathcal{S}^{\mathcal{F}}$. These may have been added to get the robot to behave in certain ways. Secondly, humans may ignore states they think are guaranteed to be reached or will never be reached.
While we can’t exactly calculate $\mathcal{S}^{\mathcal{F}}$ and $\mathcal{S}^{\mathcal{G}}$, we will proceed to show how we can generate supersets of these sets. In particular, we will calculate the set of all states not reachable under any optimal policies (denoted as $\widehat{\mathcal{S}}^{\mathcal{F}}\supseteq\mathcal{S}^{\mathcal{F}}$) and the set of all states that are reachable under every optimal policy ($\widehat{\mathcal{S}}^{\mathcal{G}}\supseteq\mathcal{S}^{\mathcal{G}}$). We can generate these sets by using modified forms of the LP formulation discussed in Equation 1. To test whether a given state $s_{i}$ is part of $\widehat{\mathcal{S}}^{\mathcal{F}}$ or not, we will look at an LP that will try to identify an optimal policy that will visit $s_{i}$. Specifically, the LP would look like:
$\begin{split}\max_{x}\sum_{s,a}x(s,a)\times r(s,a)+\alpha\times(\sum_{a}x(s_{i},a))\\[-5.0pt] \text{s.t.}\quad\forall s\in S,\sum_{a}x(s,a)=\delta(s,s_{0})+\gamma\times\sum_{s^{\prime},a^{\prime}}x(s^{\prime},a^{\prime})\times T^{H}(s^{\prime},a^{\prime},s)\\[-5.0pt] \sum_{s,a}x(s,a)\times r(s,a)=V^{*}_{s_{0}}\\[-8.0pt] \end{split}$ (2)
Where $\alpha$ is some positive valued coefficient and $V^{*}_{s_{0}}$ is the value of the optimal policy in the starting state $s_{0}$. We can calculate $V^{*}_{s_{0}}$ by solving Equation 1 on the human model with the specified reward. By adding the new term to the objective, we provide higher objective value to policies that visit the state $s_{i}$. At the same time, the constraint $\sum_{a}x(s_{0},a)=V^{*}_{s_{0}}$ ensures we only consider optimal policies. We can now formally state the following property for the above LP formulation.